### v1

# Articulated Object Manipulation using Online Axis Estimation with SAM2-Based Tracking

Abstract
Articulated object manipulation requires precise object interaction, where the object’s axis must be carefully considered. Previous research employed interactive perception for manipulating articulated objects, but typically, open-loop approaches often suffer from overlooking the interaction dynamics. To address this limitation, we present a closed-loop pipeline integrating interactive perception with online axis estimation from segmented 3D point clouds. Our method leverages any interactive perception technique as a foundation for interactive perception, inducing slight object movement to generate point cloud frames of the evolving dynamic scene. These point clouds are then segmented using Segment Anything Model 2 (SAM2), after which the moving part of the object is masked for accurate motion online axis estimation, guiding subsequent robotic actions. Our approach significantly enhances the precision and efficiency of manipulation tasks involving articulated objects. Experiments in simulated environments demonstrate that our method outperforms baseline approaches, especially in tasks that demand precise axis-based control.
I INTRODUCTION
Robotic manipulation has a wide range of applications, such as industrial automation, medical surgery, and warehouse logistics. Among various manipulation tasks, those involving articulated objects, such as doors and drawers, are particularly challenging. This is because manipulating articulated objects requires not only a deep understanding of their overall geometry but also knowledge of their part compositions and the kinematic relationships between those parts [1].
Traditional manipulation techniques, which typically rely on predefined kinematic models [2, 3] or open-loop control [4], have shown limitations in dynamically adapting to real-world interactions, especially in the presence of uncertainties. One significant drawback of open-loop approaches is their lack of feedback regulation, preventing the robot from adjusting its actions based on real-time interaction dynamics. Without continuous feedback, the system cannot adapt to
Figure 1: Our approach introduces a closed-loop pipeline integrating SAM2 to track the articulated object throughout the manipulation process captured by an RGB-D camera. The mask is subsequently used to segment out the point cloud of the articulated object, followed by explicitly estimating the joint axis using the oriented bounding box (OBB) of the motion part derived from the point cloud. The robot’s action is instructed under the guidance of the motion axis to perform further manipulation for the next loop.
changes in the object’s state, leading to inefficiencies and inaccuracies in manipulation tasks.
Recently, interactive perception has emerged as a promising approach to address these challenges [1, 5, 4, 6]. By actively interacting with the object, robots can gather real-time sensory data that provides insights into the object’s structure and kinematics. Interactive perception aims to bridge the gap between perception and action by using sensory feedback from these interactions. While this method is effective in providing information about the object’s structure and state, it lacks a critical component — the evolving dynamic interactions between the robot and the articulated object over time. This omission limits the robot’s ability to adapt its manipulation policy in real time as the object’s state changes, leading to inefficiencies in tasks that require precise control. To address this limitation, in this paper, we propose a novel closed-loop pipeline that integrates interactive perception with online refined axis estimation, providing the robot with helpful guidance for axis-aware manipulation. The basic idea is illustrated in Fig. 1. Our approach builds on interactive perception and enhances it by using the guidance of an online refined motion axis, which is derived from segmented 3D point clouds. Specifically, we use any interactive perception technique, such as RGBManip [5], leveraging it to induce slight object movements and generate dynamic 3D point cloud frames. These frames are then processed using an advanced segmentation pipeline, Grounding DINO [7] as object detector and the Segment Anything Model 2 (SAM2) [8] for object segmentation, which identifies and isolates the articulated object. By masking the moving components of the object, we can explicitly calculate the motion axis, which in turn informs the robot’s subsequent manipulation policy. In contrast to previous studies, our closed-loop integration of interactive perception with online refined axis estimation allows for more precise and adaptive control, addressing the limitations of traditional methods that operate in an open-loop fashion. By continuously updating the robot’s understanding of the object’s kinematic state, our approach ensures that the robot maintains effective control throughout the manipulation task. We validate our approach through tasks that require precise axis-based control in simulated environments, demonstrating that our method outperforms baseline approaches. Experimental results demonstrate that our pipeline significantly enhances the precision of manipulation tasks involving articulated objects, while strong generalization capability across articulated object manipulation tasks.
To summarize, the three key contributions of our paper are as follows:
  1. 1.
    Closed-loop integration of interactive perception and real-time axis estimation: By continuously refining the motion axis based on updates from 3D point clouds, our approach enables adaptive and precise control during manipulation tasks, overcoming the limitations of the open-loop methods.
  2. 2.
    Utilization of advanced segmentation models for precise articulation recognition: Our pipeline employs Grounding DINO for object detection and SAM2 for accurate segmentation, allowing the robot to isolate moving components and calculate the motion axis of moving parts of articulated objects in real-time.
  3. 3.
    Enhanced generalization and precision in articulated object manipulation: Our method significantly improves manipulation precision and generalization across diverse cabinets in door-opening and drawer-opening tasks, outperforming state-of-the-art baselines and providing consistent axis-aware manipulation.
II RELATED WORK
II-A Vision-based Robotic Manipulation
Visual perception is crucial for robotic manipulation, whose quality heavily depends on the visual modalities. Existing works on vision-based robotic manipulation have adopted different types of visual input, with various visual modalities demonstrating strong potential in different scenarios, while also exhibiting limitations specific to their applications.
RGB images, despite their richness in texture, lack depth information, which is critical for object manipulation and grasping tasks. To address this, UMPNet [9] and Grasp Proposal Networks [10] turned to depth sensors and combined RGB images with depth maps. RGB-D information provides a more direct representation of the 3D world, which, however, can be noisy and affected by transparent and reflective surfaces [11].
To circumvent the above challenge, unlike the recent approaches that switch between different visual modalities [12, 13, 14], RGBManip utilizes RGB-only visual input and directly estimates the 6D pose of objects from multi-view RGB images. It relies on a single RGB camera mounted on its end-effector as its primary visual input. As a result, during the pre-grasping phase, the camera predominantly captures only the immediate vicinity of the handle, inevitably missing crucial information about the interaction dynamics that occur throughout the manipulation process, which is critical for the robot to understand how articulated objects should be manipulated.
On the other hand, point-cloud-based methods have also demonstrated promising performance in vision-based robotic manipulation. Several studies such as Where2Act [15] and SAGCI-System [16] have integrated point cloud information to predict action affordances. Both methods are particularly well-suited for handling articulated objects, but RLAfford [10] emphasizes interaction dynamics which helps to predict manipulation policies effectively. Furthermore, Flowbot3D [17] leverages point cloud motion flows to predict how objects should be manipulated, whose strength is in handling dynamic articulated objects.
Our approach strives to integrate the strengths of RGB-based methods and point cloud-based techniques by tracking RGB inputs and segmenting the point clouds corresponding to the motion part of articulated objects. In order to address the limitation of opened-loop methods, we integrate an additional RGB-D camera into the scene, specifically aimed at capturing the manipulation process.
II-B 3D Point Cloud Segmentation Powered by 2D Foundation Models
Segmentation techniques for 2D images have reached a high level of sophistication, with models such as Segment Anything Model (SAM) [18] standing out as prominent examples. However, the limitations of segmentation methods for 3D point clouds have hindered the development of their application in robotic manipulation.
DeepLabv3 [19] has adapted 2D image segmentation techniques to the 3D point cloud domain by projecting 3D data onto 2D planes and performing segmentation within the transformed space. It exhibits robustness when handling large-scale data, but the projection from 3D point clouds onto 2D planes can lead to a loss of information, which in turn compromises segmentation accuracy.
To overcome this challenge, researchers have proposed some SAM-based methods for 3D point cloud segmentation. Point-SAM [20] extended SAM to 3D domain, which can
Figure 2: In our pipeline, an RGB-D camera captures the dynamic scene, which is induced by the slight movement from the Interactive Perception & Init-Manipulation Module. The captured scene is then processed by the Tracking & Segmentation Module, which tracks and segments the moving part of the articulated object at a 3D level. This segmented data is subsequently passed to the Axis Estimation & Manipulation Module. Here, the motion axis is explicitly calculated, providing informed guidance for the robot’s manipulation policy.
generate segmentation masks in response to point and mask prompts. For more precise segmentation such as part-level segmentation, works like PartSLIP [21] and its enhanced version PartSLIP++ [22] perform 3D point cloud segmentation by employing GLIP [23] and SAM on multiple views of a dense point cloud. All these approaches offer different solutions to the obstacle posed by the irregularity scale and complexity of 3D data, while they are also constrained by the limited viewpoints and sparse point clouds in real-world scenarios.
In practice, sparse viewpoint observations are often sufficient to meet the demands of vision-based robotic manipulation, as only a coarse segmentation of the relevant parts of the target objects is typically required. Our approach builds upon the RGBManip environment by integrating a single RGB-D camera and applying noise-robust techniques to process the point cloud, which enhances the feasibility and portability of our method in real-world scenarios.
Previous SAM-based 3D point cloud segmentation techniques have shown promising results. Here, we highlight that for dynamic scenes, leveraging Segment Anything Model 2 (SAM2) as the foundation model for 3D point cloud segmentation offers superior preservation of the scene’s dynamic information. By masking the moving parts of the target objects in the point cloud, our approach captures rich motion-related information that can better inform the manipulation policy.
II-C Axis Estimation for Articulated Objects
Axis estimation plays a critical role in the perception, comprehension and manipulation of articulated objects. Extensive research contributes various techniques to tackle the complexities of recognizing and estimating the motion constraints and joint types of articulated objects.
Individual observations are often inherently ambiguous for articulated objects. [24] highlighted the important role of interactive perception in manipulation. Subsequent research has extensively adopted interactive perception techniques to effectively mitigate the ambiguity inherent in single observations. Act the Part [25] enabled robots to learn how to interact with articulated objects, which contributes to discovering and segmenting their parts. [26] presented an online interactive perception system based on task-specific priors to extract kinematic and dynamic models of articulated objects.
However, previous studies for axis estimation are typically open-loop, which causes a loss of the interaction dynamics. To address this issue, we combine interactive perception techniques with axis estimation, introducing a novel closed-loop pipeline. With the moving components of the articulated objects segmented from each frame scene point cloud, we explicitly estimate the axis of articulation by leveraging task-based priors related to the geometric properties of articulated object motion. Our approach does not depend on high-quality point clouds or extensive manipulation, making it promising for application in complex real-world environments.
III METHOD
We introduce a novel approach for manipulating articulated objects, guided by axis estimation derived from SAM2-based tracking. As depicted in Fig. 2, our pipeline consists of 3 core modules: (1) Interactive Perception & Init-Manipulation Module, (2) Tracking & Segmentation Module and (3) Axis Estimation & Manipulation Module.
Initially, the robot applies any interactive perception methods to manipulate the articulated object, creating a slight displacement. Then, using a SAM2-based approach, the moving part of the object’s point cloud in each frame is captured, followed by further manipulation guided by the estimated axis.
III-A Interactive Perception & Init-Manipulation Module
The Interactive Perception & Init-Manipulation Module employs flexible interactive perception methods to initiate manipulation of articulated objects, producing slight displacements for further analysis. Our method incorporates the state-of-the-art RGBManip as the default implementation for this module, chosen for its recent success in interactive perception. However, the module allows for substitution with other interactive perception techniques that may be more suited to specific scenarios, providing adaptability based on application requirements. As RGBManip initiates the manipulation of articulated objects, the camera periodically records dynamic RGB-D data. The captured RGB data will be leveraged for subsequent tracking and segmentation tasks, while the corresponding depth data facilitates the reconstruction of the 3D point cloud of the evolving dynamic scene. This setup ensures precise monitoring of object motion and enhances the fidelity of axis estimation and further manipulation steps.
III-B Tracking & Segmentation Module
The Tracking & Segmentation Module plays a pivotal role in accurately tracking and segmenting the point clouds of dynamic scenes obtained from the Interactive Perception & Init-manipulation Module. Specifically, we exploit Grounding DINO on the initial RGB frame to generate an anchor box for the target object using the textual prompt. The first RGB frame serves as the starting point for SAM2’s tracking process. The anchor box, corresponding to the target object, is input into SAM2 as a box prompt, enabling it to continuously track the object and provide a mask for it throughout each moment of the dynamic manipulation process. With the object’s mask, we can extract the corresponding portion of the 3D point cloud representing the object from the dynamic scene. After a basic noise filtering step, we compute the oriented bounding box (OBB) for the entire cabinet in the initial frame. Subsequently, by subtracting this OBB from each frame’s point cloud and applying another round of noise reduction, we obtain the point cloud representing the moving parts of the articulated objects. For implementation details, the segmentation can be refined by removing the protruding handle from the cabinet’s OBB to obtain a tighter OBB of the cabinet’s body, which can lead to more precise motion-part segmentation.
III-C Axis Estimation & Manipulation Module
The Axis Estimation & Manipulation Module calculates the axis of motion based on the segmented point cloud representing the articulated object’s moving parts identified by the Tracking & Segmentation Module. Task-specific geometric priors play an essential role in achieving accurate axis estimation, i.e., for a prismatic joint, the moving parts translate along the axis, whereas for a revolute joint, the moving parts rotate around the axis. In our simulation settings, the drawer is designed to translate parallel to the xy-plane, while the door rotates about a hinge aligned with the z-axis. By leveraging these task-based priors, we are able to explicitly calculate the joint’s axis. Consider a single action of robotic manipulation, where the OBBs of the initial and final point cloud of the motion components of the articulated object are denoted as $obb_{st}$ and $obb_{ed}$, with their centers designated as $O_{st}$ and $O_{ed}$ respectively. On the one hand, for a prismatic joint, the axis pivot point is defined as the center of $obb_{st}$, and the axis direction is estimated as the direction $\overrightarrow{d}=\overrightarrow{O_{st}O_{ed}}$. On the other hand, for a revolute joint, the axis pivot point is derived by identifying the intersection of the mid perpendiculars along the longer edges in the top-down view of $obb_{st}$ and $obb_{ed}$. With the axis point $P$ established, the axis direction is ascertained by evaluating the sign of the dot product $\overrightarrow{d}\cdot\overrightarrow{t}$, where $\overrightarrow{t}=\overrightarrow{O_{st}P}\times\overrightarrow{z}$, among which $\overrightarrow{z}$ represents the positive direction of the z-axis.
We propose a closed-loop axis estimation refinement policy predicated on a straightforward observation: as the manipulation of the joint assembly increases, the point cloud of its moving component becomes progressively more amenable to accurate reconstruction. The enhancement in the fidelity of the point cloud data, in turn, furnishes the axis estimation process with inputs that are increasingly precise and dependable. For implementation, after each manipulation, we invoke the Tracking & Segmentation Module and the Axis Estimation Module to ascertain the current axis estimation. The Manipulation Module then guides the robot’s subsequent actions according to this estimation. The interactive process reiterates until the robot has executed all designated actions, with the axis estimation being continuously refined throughout the procedure.
IV EXPERIMENT
IV-A Task Settings
Our method specifically focuses on door-opening and drawer-opening tasks. We evaluate the performance of these tasks in scenarios with large ranges of door opening and drawer extension.
In each task, the articulated objects, placed with limited random position and rotation, begins in a closed state, and the robotic arm is initially positioned randomly in front of it. The arm must then accomplish the designated manipulation goal — either opening the drawer or door to a specific degree or range.
The task settings are detailed as follows:
  • Open Door: The agent needs to open the door larger than 8.6°, 10°, 20°, 30°, 40°, 45°, 50°, 55°, 60°, 65°, 70°.
  • Open Drawer: The agent needs to open the drawer larger than 10 cm, 15 cm, 20 cm, 25 cm, 30 cm, 35 cm, 40 cm, 45 cm.
To ensure fairness in comparison, all RGBManip components employ task-specific adapose as the pose estimator and heuristic pose as the controller among different tasks. All methods are compared under equivalent total step sizes, though different methods may allocate step sizes differently based on their respective policies.
For each task, we evaluate our methods compared with RGBManip and other baselines separately on RGBManip’s training set and testing set. Success rates of the first 100 experiments are used as metrics for comparison respectively.
Figure 3: Success rate of more challenging tasks for opening door.
Figure 4: Success rate of more challenging tasks for opening drawer.
TABLE I: Quantitative comparison between our method and baselines
Methods Open Door 8.6° Open Door 45° Open Drawer 15 cm Open Drawer 30cm
Train Test Train Test Train Test Train Test
DrQ-v2  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 1.8 2.5 0.8 0.8 1.9 1.0 1.4 0.5
LookCloser  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 1.5 1.25 0.8 0.8 0.8 0.0 0.0 0.0
RGBManip  [^2] Results reproduced using RGBManip under the settings outlined in Sec. IV-A. 75.0 82.0 47.0 47.0 56.0 64.0 46.0 45.0
Where2Act  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 8.0 7.0 1.8 2.0 5.9 7.5 1.1 0.6
Flowbot3D  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 19.5 20.4 6.8 6.4 27.3 25.8 16.9 11.3
UMPNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 27.1 28.1 11.0 10.9 16.6 18.8 4.4 5.6
GAPartNet  [^1] Experimental results derived from TABLE I of the paper RGBManip: Monocular Image-based Robotic Manipulation through Active Object Pose Estimation [5]. 69.5 74.5 39.4 43.6 50.6 59.3 44.6 48.6
Ours 87.0 88.0 54.0 54.0 68.0 85.0 59.0 68.0
TABLE II: More challenging tasks for opening door
Methods 20° 30° 40° 50°
Train Test Train Test Train Test Train Test
RGBManip 72.0 75.0 66.0 66.0 53.0 56.0 39.0 38.0
Ours 75.0 77.0 72.0 70.0 62.0 59.0 44.0 44.0
Methods 55° 60° 65° 70°
Train Test Train Test Train Test Train Test
RGBManip 32.0 32.0 26.0 28.0 23.0 16.0 22.0 13.0
Ours 38.0 41.0 28.0 35.0 22.0 27.0 19.0 22.0
TABLE III: More challenging tasks for opening drawer
Methods 20 cm 25 cm 35 cm 40 cm 45 cm
Train Test Train Test Train Test Train Test Train Test
RGBManip 51.0 61.0 48.0 57.0 41.0 43.0 34.0 31.0 24.0 14.0
Ours 64.0 84.0 61.0 76.0 57.0 63.0 50.0 58.0 35.0 41.0
IV-B Baselines and Quantitative Results
We benchmark our approach against seven existing techniques (including RGBManip). Below is a brief summary of each comparison method.
  • DrQ-v2 [27]: Based on reinforcement learning (RL), it takes in the robot’s state and RGB image to determine the desired 6D pose of the robot’s end-effector.
  • LookCloser [28]: A multi-perspective RL model that leverages multi-view inputs and visual transformers to amalgamate data from various angles.
  • RGBManip [5]: Utilizes RGB-only visual input to directly estimates the 6D pose of objects from multi-view images.
  • Where2Act [15]: Processes point-cloud data to estimate the best point of interaction for manipulation.
  • Flowbot3D [17]: Predicts point-wise motion, or “flow”, within the point cloud. The point with the highest motion magnitude is selected for interaction.
  • UMPNet [9]: Utilizing RGB-D images, it predicts an action point in the image and projects it into 3D space using depth data.
  • GAPartNet [29]: A pose-centric method which predicts the pose of an object’s part from point-cloud data.
Quantitative results of basic tasks are summarized in Table. I, where both our method and RGBManip almost outperform other baseline approaches while Ours consistently surpasses RGBManip in basic tasks.
Granular results for more challenging door-opening tasks are presented in Table. II, with the corresponding line chart showing success rates with regard to door-opening angles depicted in Fig. 3.
Fine-grained results for more challenging drawer-opening tasks are provided in Table. III, with the corresponding line chart illustrating success rates at various drawer-opening distances shown in Fig. 4. Experimental results show that Ours consistently outperforms RGBManip with a significant enhancement in success rates.
V Analysis
The analysis section delves into the specific experimental results to elucidate the practical advantages of our online axis estimation approach in the context of articulated object manipulation. Our method’s performance is grounded in the empirical data obtained from the experiments, which are analyzed below to provide a detailed understanding of the improvements achieved.
V-A Online Axis Estimation vs. Traditional Methods
Our experiments clearly demonstrate the superiority of our online axis estimation approach over traditional methods, especially in tasks that demand precise control such as door and drawer opening. As illustrated in Table I, our method achieves an impressive 87.0% success rate in the training set and 88.0% in the test set for the “Open Door” task, significantly outperforming RGBManip, which records 75.0% and 82.0% respectively. This notable enhancement stems directly from the continuous refinement of the axis estimation, empowering the robot to adapt its actions in response to the most current interaction dynamics.
Furthermore, in the more demanding versions of these tasks, depicted in Figures 3 and 4, our method’s reliance on ongoing online axis estimation reveals enhanced robustness in handling large-amplitude movements of articulated objects. The experimental data consistently show that our method sustains higher success rates as the complexity of tasks escalates, whether it involves opening doors to wider angles or drawers to greater extents. The online axis estimation process is pivotal in this regard, enabling our system to dynamically adjust to real-time interaction feedback and ensuring precise control over articulated objects. The continuous axis refinement, aligned with the object’s changing state, is essential for the manipulation’s accuracy and efficiency, particularly in tasks with substantial state variations. The real-time feedback loop is crucial, enabling the robot to make swift and accurate operational adjustments that accommodate the subtleties of the interaction, thereby achieving a higher success rate in task completion. This robust performance underscores the strength of online axis estimation in providing reliable and responsive control in sophisticated robotic manipulation scenarios.
V-B Consistency Across Various Manipulation Scenarios
Our method’s consistent performance across a range of manipulation scenarios highlights its robustness and versatility. The success rates in both door and drawer opening tasks, as depicted in Tables II and III, consistently show higher rates for our method, indicating that the online axis estimation is effective regardless of the specific manipulation task. This consistency is a significant advantage over methods that may perform well in one scenario but falter in others.
VI CONCLUSION
In this research, we introduced a novel approach to articulated object manipulation that employs online axis estimation integrated with SAM2-based tracking. Our method utilizes segmented 3D point cloud data and motion axis estimation to enhance the precision and control of robotic actions. The experimental results suggest that our method shows promising improvements in handling large manipulation movements, indicating the potential utility of axis-based guidance in robotic manipulation tasks.
The online axis estimation technique has demonstrated beneficial in dynamically adjusting to the real-time interaction feedback, thereby enabling the robot to maintain better control over articulated objects. While our method has shown enhancements over traditional approaches, particularly in complex manipulation scenarios, it is important to acknowledge that there is still room for further refinement and optimization.
Our findings indicate that the online axis estimation process contributes positively to the accuracy and efficiency of robotic manipulation. However, this study represents an initial step, and future work is necessary to fully explore the long-term reliability and scalability of this approach across a broader spectrum of tasks and environments.